A canoeist paddled upstream at 10 meters per minute, turned around, and drifted downstream at 15 meters per minute. If the distance traveled in each direction was the same, and the time spent turning the canoe around was negligible, what was the canoeist's average speed over the course of the journey, in meters per minute? anyone?
Answer by stanbon(48558)

A canoeist paddled upstream at 10 meters per minute, turned around, and drifted downstream at 15 meters per minute. If the distance traveled in each direction was the same, and the time spent turning the canoe around was negligible, what was the canoeist's average speed over the course of the journey, in meters per minute? anyone?
-----------------
Upstream DATA:
rate 10 m/min ; distance = x meters ; time = d/r = x/10 minutes
-------------------------
Downstream DATA:
rate = 15 m/min ; distance = x meters ; time = d/r = x/15 min
--------------------------
Roundtrip DATA:
distance = 2x meters ; time = (x/10)+(x/15) = (3x+2x)/30 = x/6 minutes
----------------
Equation:
Avr. Rate = total distance / total time = 2x/(x/6) = 12 meters per minute
================================
Cheers,
Stan H.